OpenStudy (anonymous):

Solve the equation by completing the square: x2 + 2x = 3 *NOTE: Applicable to this and all other questions-- For any solution that involves two answers, type one answer separated by a comma and then the other answer.

OpenStudy (anonymous):

is it -3,1

OpenStudy (anonymous):

\[x^2 +2x = 3\] \[x^2 +2x - 3 =0\] \[x^2 +3x -x - 3 =0\] solve for factors

OpenStudy (anonymous):

um how

OpenStudy (anonymous):

\[x(x +3) -1(x + 3) =0\] \[(x +3)(x -1) =0\]

OpenStudy (anonymous):

so i mutiply 3 and -1

OpenStudy (anonymous):

no x+3 = 0 and x-1 = 0 are your answers that is x = -3,1